OpenStudy (anonymous):

John has taken out a loan for college. He started paying off the loan with a first payment of $100. Each month he pays, he wants to pay back 1.1 times as the amount he paid the month before. Explain to John how to represent his first 20 payments in sigma notation. Then explain how to find the sum of his first 20 payments, using complete sentences. Explain why this series is convergent or divergent.

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

@Safa102 can you help me plz?

OpenStudy (safa102):

ok so i don't know what sigma notation is but i'm pretty sure what u have to do is, do 20 divided by 1.1 than get your answer than subtract that by 20..i'm not sure but this is what i got. i hope this helps :)

OpenStudy (anonymous):

Okay thanks :)

OpenStudy (safa102):

np! :)

OpenStudy (anonymous):

@Kainui plz help?

OpenStudy (anonymous):

@hartnn can you help me plz?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

This is a geometric series, so you'll use the nth term of the geometric sequence formula, which is \[\Large a_n = a*(r)^{n-1}\] look familiar?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

what would the value of 'a' and 'r' be?

OpenStudy (anonymous):

A = 20 R = 1.1?

jimthompson5910 (jim_thompson5910):

a = 20 is not correct

OpenStudy (anonymous):

So would a = 100?

jimthompson5910 (jim_thompson5910):

yes that's the starting amount

OpenStudy (anonymous):

So it will be 100 * (1.1)^n-1 Where n = 20

jimthompson5910 (jim_thompson5910):

you're close but not quite there

jimthompson5910 (jim_thompson5910):

\[\Large \sum_{k=0}^{n}\left[a*(r)^{n-1}\right] = a*\frac{1-r^n}{1-r}\] \[\Large \sum_{k=0}^{n}\left[100*(1.1)^{n-1}\right] = 100*\frac{1-(1.1)^n}{1-1.1}\]

OpenStudy (anonymous):

Sorry I don't get this. Can you please explain?

jimthompson5910 (jim_thompson5910):

sorry I made a typo, let me fix

jimthompson5910 (jim_thompson5910):

\[\Large \sum_{k=0}^{n}\left[a*(r)^{k-1}\right] = a*\frac{1-r^n}{1-r}\] \[\Large \sum_{k=0}^{n}\left[100*(1.1)^{k-1}\right] = 100*\frac{1-(1.1)^n}{1-1.1}\]

jimthompson5910 (jim_thompson5910):

are you familiar with sigma notation? and what it represents?

OpenStudy (anonymous):

Not really

jimthompson5910 (jim_thompson5910):

sigma is used for summations so adding numbers up but instead of saying something like 1+2+3+4+5+6+7+8+9+10 which is a lot to write down, we can condense all that into \[\Large \sum_{k=1}^{10}k\]

jimthompson5910 (jim_thompson5910):

so when they're asking for sigma notation for the payments, they are asking you to add up all of the payments. You could take the long way of adding up each one by one or you can use the shortcut formula

OpenStudy (anonymous):

So since he payed 100 first and wants to pay 1.1 times the amount in his first 20 payments it will be 1.1*100

jimthompson5910 (jim_thompson5910):

does my example where I wrote out `1+2+3+4+5+6+7+8+9+10` make sense?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

the first payment is $100, this is given what is the next payment equal to?

OpenStudy (anonymous):

110

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

how about the next payment?

OpenStudy (anonymous):

121

jimthompson5910 (jim_thompson5910):

yes, so we're going to have 100+110+121+... all the way up to 20 payments

OpenStudy (safa102):

:O your still at it...wow

jimthompson5910 (jim_thompson5910):

now you could find all 20 values and add them all up, or you can take the shortcut which is what I recommend so.... \[\Large a*\frac{1-r^n}{1-r} = 100*\frac{1-(1.1)^{20}}{1-1.1} = ??\] a = 100 is the first payment r = 1.1 is the multiplier n = 20 is the number of payments

OpenStudy (anonymous):

5727.49

jimthompson5910 (jim_thompson5910):

I'm getting `5,727.4999493256` which rounds to `5,727.50`

OpenStudy (anonymous):

Ok thanks!

OpenStudy (anonymous):

I get it now :)

jimthompson5910 (jim_thompson5910):

you're welcome